OpenStudy (anonymous):

How is the graph of y = –x2 – 1 different from the graph of y = –x2?

OpenStudy (anonymous):

you should notice that y=-x^2 is on unit up from y=-x^2 -1, in other words y=-x^2-1 intercepts the y axis at -1 where y=-x^2 intercepts at the origin

OpenStudy (anonymous):

ok thank you so much